OpenStudy (anonymous):

can anyone help me solve the lim of lnx/x-1 as x goes to 1

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}\ln x \div x-1\]

OpenStudy (lgbasallote):

\[\lim_{x\rightarrow 1} (\frac{\ln x}{x-1})?\]

OpenStudy (lgbasallote):

coincidentally...if you substitute x = 1 right now..you'll get 0/0 <--you can use l'hospital

OpenStudy (lgbasallote):

pls tell me you know l'hospital....

OpenStudy (lgbasallote):

i hate limits that dont use l'hospital lol

OpenStudy (anonymous):

yes I know

OpenStudy (anonymous):

and yes that is the right equation

OpenStudy (lgbasallote):

ahh wonderful!

OpenStudy (lgbasallote):

take the derivative of numerator and denominator \[\frac{\frac{1}{x}}{1} \implies \frac 1x\] now sub in \[\lim_{x \rightarrow 1}\frac{1}{1} \implies 1\]

OpenStudy (lgbasallote):

do you get that?

OpenStudy (lgbasallote):

@Zarkon feel free to point out if i made a mistake...limits is not my field of expertise i admit

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

nice!

OpenStudy (anonymous):

trust me I'm just learning this stuff and I only have 3 weeks to get this Calculus class over.

OpenStudy (lgbasallote):

l'hospital is the only topic i remember from limits lol

OpenStudy (anonymous):

lol

OpenStudy (lgbasallote):

i suggest you practice your derivatives though...i believe that makes up 3/4 of the calculus I curriculum so if you master derivatives..you're good...and it'll make it easier for you to do l'hospital

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

can someone help me slove this attachment below.